# Design and deployment of a fast neural network for measuring the properties of muons originating from displaced vertices in the CMS Endcap Muon Track Finder

### v1

@@ 2 Design and Hardware Implementation
Therefore an NN-based solution is chosen to be implemented in Run 3 EMTF system that targets measurement of $p_{T}$ and $d_{0}$ of displaced muons. Two separate NNs are developed in software to target $p_{T}$ and $d_{0}$ separately. Each of the NN models have three dense layers, and equally split half of the total nodes of the complete model. The two models are then stitched together to form a single NN model before being implemented into the FPGA. The models are trained on simulations of LLPs that decay to muons made with standard CMS software emulators.
@@
Final implementation of the NN model fits into a latency of 83 ns. After the NN implementation the LUT usage in the FPGA is 76% (from 74%) and the DSP usage is 14% (from 2%). The validation of the model shows good agreement between $p_{T}$ and $d_{0}$ measured by the NN model and the true values. The underestimation of $p_{T}$ at higher true $p_{T}$ is not crucial since the current displaced muon trigger designs target 5-10 GeV range which has good accuracy.

### v2

@@ 2 Design and Hardware Implementation
Therefore an NN-based solution is chosen to be implemented in Run 3 EMTF system that targets measurement of $p_{T}$ and $d_{0}$ of displaced muons. Two separate NNs are developed in software to target $p_{T}$ and $d_{0}$ separately. Each of the NN models have three dense layers, and equally split half of the total nodes of the complete model as seen in Fig. 1. The two models are then stitched together to form a single NN model before being implemented into the FPGA. The models are trained on simulations of LLPs that decay to muons made with standard CMS software emulators.
@@
Final implementation of the NN model fits into a latency of 83 ns. After the NN implementation the LUT usage in the FPGA is 76% (from 74%) and the DSP usage is 14% (from 2%). The validation of the model shows good agreement between $p_{T}$ and $d_{0}$ measured by the NN model and the true values as shown in Fig. 2. The underestimation of $p_{T}$ at higher true $p_{T}$ is not crucial since the current displaced muon trigger designs target 5-10 GeV range which has good accuracy.